Problem #6-15 PTS many identical particles

Consider a one-dimensional infinite square well of width w= 1 cm with free electrons in it.

if the Fermi energy is 2ev, what is the number of electrons inside the well?



1
Expert's answer
2021-06-29T07:03:02-0400

Width of the well w = 1 cm = 0.01 m

Fermi energy "E_F= 2\;eV"

Fermi energy

"E_F = \frac{h^2}{2m_e}(\frac{3π^2N}{V})^{2/3}"

me = rest mass

V = volume of the system

V=0.01

"N = \frac{E_F^{3/2} \times 2m_eV}{h^2 \times 3π^2} \\

= \frac{2^{3/2} \times 2 \times 9.1 \times 10^{-31} \times 0.01}{(1.03 \times 10^{-34})^2 \times 3π^2} = 1.73 \times 10^{35}"

Answer: number of electrons "1.73 \times 10^{35}"